# A Comprehensive Evaluation of Large Language Models on Temporal Event Forecasting

### v1

@@ 3. Methods
We systematically designed a series of LLM-based methods for event forecasting, categorized into three streams: pure graph-based, pure text-based, and mixed graph-text methods. Notably, we only introduce methods that employ LLM, while non-LLM baselines are described in Section 4.1. Additionally, we developed two methods for constructing event history to assist LLM reasoning in temporal event forecasting, as illustrated in Figure 2. These methods include rule-based history and retrieved history.
3.1. Graph-only Methods
@@
Graph-only methods utilize discrete event graphs as inputs, leveraging the inherent domain knowledge of LLMs to reason over historical events and subsequently make forecasts. To capitalize on contextual information and historical developments related to current query event, we develop two methods for extracting pertinent historical events from event graphs. The input format of the graph is shown in Table 3
@@
Rule-based History. Typically, the causes of an event can be categorized into internal and external reasons. Internal reasons pertain to factors inherent to the subject of the current event, while external reasons originate from the surrounding environment. Similarly, in temporal event forecasting, both internal and external reasons must be considered. On the one hand, we reconstruct the event graph $\mathbf{G}_{<t}^{q}=\left\{G_{0}^{q},G_{1}^{q},\ldots,G_{t-1}^{q}\right\}$ from global view to provide historical patterns and trends of the current event, where $G_{t^{\prime}}^{q}$ represents historical events at timestamp $t^{\prime}$ with the same subject as the query. On the other hand, we extract the graph from a local view to offer background information or specific situations for reasoning of LLMs. Formally, complex event graph $\mathbf{G}_{<t}^{c}=\left\{G_{0}^{c},G_{1}^{c},\ldots,G_{t-1}^{c}\right\}$, where $G_{t^{\prime}}^{c}$ represents historical events at timestamp $t^{\prime}$ within the same complex event. It is noted that due to the unique background information of the current event at specific stages, we only retain historical events from the most recent two days in relative time.
Retrieved History. Historical event graphs serve as rich sources of contextual information. Despite this richness, the intricate nature of the temporal event graph means that the existing graph still contains substantial noise. Therefore, unlike rule-based history, we retrieve relevant entities from the graph to obtain corresponding historical events. Specifically, we first extract entities from the historical events in global view with the same subject as the query. LLMs are then employed to eliminate irrelevant entities. The prompt of graph retrieval is implemented in Appendix  A.1. After identifying the pertinent entities $\mathbf{E}_{<t}=\left\{E_{0},E_{1},\ldots,E_{t-1}\right\}$, we reconstruct the history based on the events involving these entities.
3.2. Text-only Methods
@@
Text-only methods utilize raw textual data for forecasting. Table 3 illustrates the input format of the text. Each atomic event originates from a specific news document. Similar to graph-only methods, relevant event history is obtained through rule or retrieval.
@@
Rule-based History. Generally, news documents provide more granular contextual information for temporal event forecasting. Exploring textual information relevant to the current event becomes a critical step. Firstly, we identify the events by the historical events graph from the global view and local view and find the corresponding news documents set $\mathbf{D}_{<t}^{p}=\left\{D_{0}^{p},D_{1}^{p},\ldots,D_{t-1}^{p}\right\}$ through the relationships between documents and events. Considering that news documents usually contain a large amount of irrelevant description about the query event, we utilize LLM to summarize the news documents set, providing concise summaries $\mathbf{S}_{<t}^{p}=\left\{S_{0}^{p},S_{1}^{p},\ldots,S_{t-1}^{p}\right\}$ that retain only the core thematic information of the news. The prompt of news text summarization is implemented in Appendix  A.1. Unlike methods that rely solely on temporal event graphs, the text-only method aims to predict the missing objects based on historical news summaries.
@@
Retrieved History. Inspired by the recent research of RAG (Sun et al., 2023; Liao et al., 2023), we utilize the embeddings from LLM to directly retrieve relevant news text from a set of historical news documents $\mathbf{D}_{<t}^{c}=\left\{D_{0}^{c},D_{1}^{c},\ldots,D_{t-1}^{c}\right\}$ in the same complex event. Following this, we sort the retrieved news texts by time and filter out historical texts that are excessively outdated in terms of the event timestamp. Similarly to rule-based history, we employ LLM to generate concise summaries $\mathbf{S}_{<t}^{c}=\left\{S_{0}^{c},S_{1}^{c},\ldots,S_{t-1}^{c}\right\}$ for the filtered news text set, serving as contextual information to predict missing object in the current event.
3.3. Graph-and-Text (Mixed) Methods
@@
Graph-and-text methods leverage both graph and text data as input. The input format of mixed methods is shown in Table  3. Differing from the above two methods, the graph-and-text method combines event graphs and news documents, with the event graph supplying structured information and the news documents providing fine-grained background information. The two specific implementations of this method are as follows:
@@
Rule-based History. The rule-based history with graph-and-text integrates graph-only history with text-only history to form a comprehensive graph-and-text history. The input of LLM includes not only historical events but also their corresponding news summaries. The graph history supplies pertinent temporal event graph structure details. Concurrently, the text history furnishes detailed background information in a granular manner.
@@
Retrieved History. Rather than directly concatenating graph-only history with text-only history, the retrieval-based method initially obtains the relevant entity set by the historical event graph. Compared to the text-only retrieval method, the mixed retrieval method retrieves news text not only based on the subject of the query event but also the relevant entity set, thus augmenting the text exploration with a certain level of contextual information. The retrieved news text, in turn, better assists LLM in conducting structured reasoning.
4. Experiments
@@
4.1. Experimental Settings
To evaluate various methods, we conduct experiments on our proposed dataset MidEast-TE-m, as described in Section 2.2. We follow previous approaches and use Accuracy (Acc) as the evaluation metric. For non-zero shot methods, we train the models on the training set, choose the best-performing model based on the validation set, and finally obtain the performance by testing the model on the testing set. For zero-shot methods, we directly test them on the testing set and record the performance.
4.1.1. Compared Methods
@@
  • ConvTransE (Shang et al., 2019): it is a static knowledge graph (KG) representation learning method, which uses both a convolutional neural network and a translational operation to capture the patterns within triplets.
  • RGCN (Schlichtkrull et al., 2018): this is also a static KG representation learning model, which leverages graph convolutional neural network to model various relations among entities.
  • RE-GCN (Li et al., 2021b): REGCN is a SOTA method for TKG, which leverages GNN and RNN to capture both relational and temporal patterns, respectively.
  • CMF (Deng et al., 2021): CMF proposes a context-based feature fusion method, which integrate multilevel features including event frequency, news documents and event graphs.
  • LoGo (Ma et al., 2023b): LoGo is the SOTA method for the temporal complex event (TCE), which employs two RT-GNN modules to model both local (within complex event) and global (across all complex events) contexts. To be noted, only LoGo leverage the complex event information, while the above three methods do not use it.
  • ForecastQA (Jin et al., 2021): ForecastQA simulate the the forecasting scenario on temporal news documents and design a method based on pretrain language models to make a forecasting judgement with the retrieved documents.
  • CoH (Luo et al., 2024): It is a LLM-based method that construct the event history by designed rules and employ fine-tuning to understand textual graph information.
  • GenTKG (Liao et al., 2023): this is a retrieval-augmented generation framework that incorporates temporal logical rules into the retrieval process. Additionally, fine-tuning is employed to adapt LLM for the task of temporal event forecasting.
Table 4. Performance (Accuracy) comparison of LLM-based methods and non-LLM methods. ”N/A” stands for ”Not Applicable”.
@@
Non-LLM ConvTransE (Shang et al., 2019) GNN 0.3737 N/A N/A 0.7327 N/A N/A
RGCN (Schlichtkrull et al., 2018) 0.3777 N/A N/A 0.7203 N/A N/A
RE-GCN (Li et al., 2021b) 0.3879 N/A N/A 0.7333 N/A N/A
CMF (Deng et al., 2021) N/A N/A 0.3783 N/A N/A 0.7265
LoGo (Ma et al., 2023b) 0.3969 N/A N/A 0.7406 N/A N/A
ForecastQA (Jin et al., 2021) Bert N/A 0.3901 N/A N/A 0.7389 N/A
@@
Fine-tune CoH (Luo et al., 2024) llama2-7b 0.4856 N/A N/A 0.7763 N/A N/A
GenTKG (Liao et al., 2023) 0.4785 N/A N/A 0.7907 N/A N/A
@@ 4.1.2. Implementation Details
For the rule-based method, we generate corresponding summaries for all documents in advance with Vicuna-7b. Notably, we further apply fine-tuning (instruction tuning with QLoRA (Dettmers et al., 2023)) to these open-source LLMs. Since the retrieved history directly retrieves news text, we utilize gpt-3.5-turbo to generate summaries for the retrieved news text set and keep the retrieval models fixed without tuning. The retrieval models include three different retrievers: BM25 (Robertson et al., 2009), Contriever (Izacard et al., 2021), and LlamaIndex (Liu, 2022).To ensure the fairness of the experiment, for all the baselines and LLM-based methods, we search historical lengths from $\{3,7,15,30,90\}$. Moreover, considering the limitation of the context window, the maximum number of events and text in history is set to 20 and 5 respectively.
4.2. Performance Comparison (RQ1)
@@
4.2.1. Performance w.r.t. Various Inputs and Outputs.
@@
Table 4 demonstrates the overall performance comparison on the task of object entity prediction and relation prediction, respectively. Clearly, we have the following observations. First, the LLM-based method proposed in this paper achieves the best performance in the fine-tune setting. The preliminary results imply that there is great potential of fine-tuning LLMs for event forecasting, and more sophisticated fine-tuning is worth further exploration in the future. Compared to the other fine-tuned LLM-based methods such as CoH and GenTKG, the introduction of complex events and news text provides more fine-grained information for temporal event forecasting. Moreover, enhanced with the retrieval modules, LLM-based methods in zero-shot setting can achieve not bad performance for both tasks of tail entity and relation prediction. Nevertheless, they still under-perform conventional non-LLM-based methods on all three formats of inputs, aka. Graph-only, Text-only, and Mixed, demonstrating that non-LLM-based methods remain competitive. Second, surprisingly, incorporating raw text into the input does not demonstrate positive effects in zero-shot setting. The performance on the text-only setting is obviously worse than the graph-only and mixed settings. This may be because the raw textual inputs are longer than the graph-only inputs, resulting in increasing difficulty in reasoning. Third, comparing the two forecasting tasks, the absolute performance of relation prediction is higher than object entity prediction by an obvious margin, indicating that forecasting entities are more challenging than relations in the domain of international cooperation and conflict events. The possible reasons are from two aspects. First, the entity set is much larger than the relation set, therefore, predicting entity is more difficult than relation given a larger candidate set. Second, when predicting relations, the input question includes both the subject and object entity, while the subject entity and relation are the input when predicting object entity. We deem that entities retain more information than relation in event forecasting, thus the question of relation prediction could offer more information and is easier.
4.2.2. Performance w.r.t. Backbone LLMs
@@
The capability of the backbone LLM plays a crucial role in various tasks. In addition to the commercial proprietary LLMs, we are also interested in the performance of open-source LLMs. These open-source LLMs can be adapted, re-configured, and more importantly, safer than commercial LLMs due to privacy issues. Hence, we choose two popular open-source LLMs, i.e., LLaMA2-7b and Vicuna-7b, to test. The results are presented in Table 5. Clearly, both LLaMA2-7b and Vicuna-7b perform significantly worse than gpt-3.5-turbo on the zero-shot setting, owing to the inherent capacity gap between open-source smaller models and commercial powerful models. Nevertheless, Under the fine-tuning setting, all the open-source models exhibit significant performance gains. They are not only better than the zero-shot version counterparts but also outperform gpt-3.5-turbo by a large margin. This phenomenon indicates that supervised instruction fine-tuning is crucial for enhancing the ability of LLM to learn temporal relational patterns. Additionally, the performance of Vicuna-7b outperforms LLaMA2-7b, implying that advanced LLM capabilities are more effective at capturing evolutionary patterns in historical events.
@@
Table 5. Performance comparison w.r.t. different backbone LLMs and fine-tuning. ”N/A” means ”Not Applicable”.
Model LLaMA2-7b Vicuna-7b gpt-3.5-turbo
Zero-shot without history 0.1976 0.1812 0.2746
@@
rule-based history graph 0.2225 0.2027 0.3290
rule-based history text 0.2639 0.2384 0.3233
rule-based history Mix 0.2157 0.2524 0.3154
Fine-tune rule-based history graph 0.5057 0.5271 N/A
rule-based history text 0.4892 0.5266 N/A
rule-based history Mix 0.4977 0.5798 N/A
4.3. Effects of RAG (RQ2)

### v2

@@ 2 Methods
We systematically designed a series of LLM-based methods for event forecasting, categorized into three streams: graph-only, text-only, and graph-text (mixed) methods. Notably, we focus on LLM-based methods, while non-LLM baselines are described in Section 3.1. Additionally, we developed two approaches for constructing event history to assist LLM reasoning in temporal event forecasting, as illustrated in Figure 3. These approaches include rule-based history and retrieved history.
Figure 3: Illustration of rule-based history and retrieved history. The rule-based history is constructed by a set of predefined rules. In contrast, the retrieved history dynamically searches context from the temporal knowledge graph or news documents according to the current query.
2.1 Graph-only Methods
@@
Graph-only methods utilize discrete temporal knowledge graphs as inputs, leveraging the inherent domain knowledge of LLMs to reason over historical events and make forecasts. To enhance contextual understanding, we developed two methods for constructing event history relevant to the query event.
@@
Rule-based History. Inspired by the previous work [6], we construct event history from local and global views. In the global view, we reconstruct the knowledge graph $\mathbf{G}_{<t}^{q}=\left\{G_{0}^{q},G_{1}^{q},\ldots,G_{t-1}^{q}\right\}$ based on the connections of the query subject $s$ within the original temporal knowledge graph (TKG), where $G_{t^{\prime}}^{q}=\left\{(s,r_{i},o_{j},c_{k},t^{\prime})|(s,r_{i},o_{j},c_{k},t^{\prime})\in G_{t^{\prime}}\right\}$ represents historical events at timestamp $t^{\prime}$ with the same subject as the query. In the local view, we extract a subgraph to provide fine-grained contextual information for LLM reasoning. Formally, complex knowledge graph $\mathbf{G}_{<t}^{c}=\left\{G_{0}^{c},G_{1}^{c},\ldots,G_{t-1}^{c}\right\}$, where $G_{t^{\prime}}^{c}=\left\{(s_{i},r_{j},o_{k},c,t^{\prime})|(s_{i},r_{j},o_{k},c,t^{\prime})\in G_{t^{\prime}}\right\}$ represents historical events at timestamp $t^{\prime}$ in the same complex event $c$. Due to length constraints on event history, we retain only the most recent Top-K events for $\mathbf{G}_{<t}^{q}$ and events from the past two days for $\mathbf{G}_{<t}^{c}$ to construct the event history.
Retrieved History. Unlike rule-based history, retrieved history typically leverages the powerful capabilities of LLMs to identify relevant contextual information. Firstly, we selected the historical events from the past days within $\mathbf{G}_{<t}^{c}$ to form an entity candidate set, which represents the potential actors of the query event. This candidate set is then input to the LLM to filter out irrelevant entities. The instruction of graph retrieval is detailed in Table  3. Once the relevant entities, denoted as $\mathbf{E}_{<t}^{c}=\left\{E_{0},E_{1},\ldots,E_{t-1}\right\}$, are identified, we extract the historical events that containing $e^{c}\in\mathbf{E}_{<t}^{c}$ to construct the event history.
Table 3: Instruction of graph retrieval.
2.2 Text-only Methods
@@
Text-only methods utilize raw textual data for forecasting. Each atomic event originates from a specific news document. Similar to graph-only methods, relevant event history is obtained through rule or retrieval.
@@
Rule-based History. Generally, news documents provide more detailed contextual information for temporal event forecasting, making the exploration of textual information related to the current event a critical step. Following the graph-only method, we identify the historical events from $\mathbf{G}_{<t}^{q}$ and $\mathbf{G}_{<t}^{c}$ and construct the news documents set $\mathbf{D}_{<t}^{p}=\left\{D_{0}^{p},D_{1}^{p},\ldots,D_{t-1}^{p}\right\}$, from which these historical events can be extracted. Given that news documents often contain a substantial amount of irrelevant information regarding the query event, we employ LLMs to summarize these documents, producing concise summaries $\mathbf{H}_{<t}^{p}=\left\{H_{0}^{p},H_{1}^{p},\ldots,H_{t-1}^{p}\right\}$ that retain only the core thematic information. Compared to methods that rely solely on temporal knowledge graphs, the text-only method focuses on predicting the missing objects based on historical news summaries.
@@
Retrieved History. Inspired by the recent studies of RAG [26, 24], we construct the news document set $\mathbf{D}_{<t}^{c}=\left\{D_{t-h}^{c},\ldots,D_{t-2}^{c},D_{t-1}^{c}\right\}$ from the past $h$ days within the same complex event as the query event. To obtain the fine-grained textual information, these news documents are divided into smaller chunks, which are then encoded into vector representations using an embedding model. The query subject is utilized to compute the similarity scores and find relevant news text. The retrieved news texts are subsequently sorted by time and similarity scores, with historical texts that are excessively outdated relative to the event timestamp being filtered out. Similarly to rule-based history, we employ LLM to generate concise summaries $\mathbf{H}_{<t}^{c}=\left\{H_{0}^{c},H_{1}^{c},\ldots,H_{t-1}^{c}\right\}$ for the filtered news text set, serving as contextual information to predict missing object in the current event.
2.3 Graph-and-Text (Mixed) Methods
@@
Graph-and-text methods leverage both graph and text data as input. Differing from the above two methods, the graph-and-text method combines TKG and news documents, with the TKG supplying structured information and the news documents providing fine-grained background information. The two specific implementations of this method are as follows:
@@
Rule-based History. The rule-based history with graph-and-text integrates graph-only history with text-only history to form a comprehensive graph-and-text history. The input to the LLM includes not only historical events but also their corresponding news summaries. The graph history provides pertinent details about the temporal knowledge graph’s structure, while the text history offers detailed background information in a more granular manner.
@@
Retrieved History. Rather than directly concatenating graph-only history with text-only history, the mixed retrieval method retrieves news text not only based on the subject of the query event but also the relevant entity set $\mathbf{E}_{<t}^{c}$, thereby enriching the text exploration with structural graph information. The retrieved news text, in turn, better assists LLM in conducting structured reasoning. To address the limitation of context length, we also transform the retrieved news text into summaries by LLMs.
Table 4: Instruction of reasoning.
2.4 Prompt Design
To effectively guide LLMs in performing temporal event forecasting, we design a specific prompt for this task, which consists of an instruction and an input component. As shown in Table 4, the instruction includes the definitions of atomic and complex events, the meanings of the identifiers used in the input, and an overview of the temporal event forecasting task. Notably, red identifiers represent those used for the rule-based history, while blue identifiers denote those used for the retrieved history. As depicted in Figure 3 , the input comprises the query, historical events, and the options. The format for the input of historical events is detailed in Table 5.
Table 5: Input format of historical events.
Input Format Historical Events
Graph-only (Pakistan Foreign Ministry, Cooperate economically, Egyptian counterpart officials, 2190);
Text-only [Date]2190: Pakistani Foreign Minister Shah Mahmood Qureshi expressed great compatibility…
Mixed [Date]2189: Foreign Minister Shah Mahmood Qureshi visited Egypt… (Shah Mahmood Qureshi, Make optimistic comment, Pakistan);
2.5 Fine-tuning
Previous work has demonstrated the effectiveness of fine-tuning LLMs on temporal event forecasting[24, 25]. In order to comprehensively evaluate the impact of fine-tuning, we adopt the Low-Rank Adaptation (LoRA)[33] for open-source LLMs. The core idea of LoRA is to decompose the weight matrices $\mathbf{W}$ of a pre-trained model into a sum of a low-rank matrix and a residual matrix, typically expressed as $\mathbf{W}^{\prime}=\mathbf{W}+\bigtriangleup\mathbf{W}$, where where $\mathbf{W}^{\prime}$ is the updated weight matrix after fine-tuning, and $\bigtriangleup\mathbf{W}=\mathbf{A}\mathbf{B}$ is the low-rank adaptation matrix; $\mathbf{A}$ and $\mathbf{B}$ are two smaller matrices that allow the model to adapt to new tasks. Due to limitations in training data and time constraints, we restricted fine-tuning to the rule-based history. Specifically, we integrated the rule-based history into the designed prompt and computed the cross-entropy loss between the predicted object and the actual answers for optimization.
3 Experiments
@@
3.1 Experimental Settings
To evaluate various methods, we conduct experiments on our proposed dataset MidEast-TE-m, as described in Section 1.2. Following previous studies, we adopt Accuracy (Acc) as the evaluation metric. For non-zero shot methods, we train each model on the training set, select the best-performing model based on the validation set, and then evaluate it on the test set. In contrast, for zero-shot approaches, we directly evaluate on the test set without any additional training.
3.1.1 Compared Methods
@@
  • 1.
    ConvTransE [8]: it is a static knowledge graph (KG) representation learning method, which uses both a convolutional neural network and a translational operation to capture the patterns within triplets.
  • 2.
    RGCN [9]: this is also a static KG representation learning model, which leverages graph convolutional neural network to model various relations among entities.
  • 3.
    RE-GCN [13]: REGCN is a SOTA method for TKG, which leverages GNN and RNN to capture both relational and temporal patterns, respectively.
  • 4.
    CMF [18]: CMF proposes a context-based feature fusion method, which integrate multilevel features including event frequency, news documents and event graphs.
  • 5.
    LoGo [6]: LoGo is the SOTA method for the temporal complex event (TCE), which employs two RT-GNN modules to model both local (within complex event) and global (across all complex events) contexts. To be noted, only LoGo leverage the complex event information, while the above three methods do not use it.
  • 6.
    ForecastQA [17]: ForecastQA simulate the the forecasting scenario on temporal news documents and design a method based on pretrain language models to make a forecasting judgement with the retrieved documents.
  • 7.
    CoH [25]: It is a LLM-based method that construct the event history by designed rules and employ fine-tuning to understand textual graph information.
  • 8.
    GenTKG [24]: this is a retrieval-augmented generation framework that incorporates temporal logical rules into the retrieval process. Additionally, fine-tuning is employed to adapt LLM for the task of temporal event forecasting.
Table 6: Performance (Accuracy) comparison of LLM-based methods and non-LLM methods. ”N/A” stands for ”Not Applicable”.
@@
Non-LLM ConvTransE [8] GNN 0.3737 N/A N/A 0.7327 N/A N/A
RGCN [9] 0.3777 N/A N/A 0.7203 N/A N/A
RE-GCN [13] 0.3879 N/A N/A 0.7333 N/A N/A
CMF [18] N/A N/A 0.3783 N/A N/A 0.7265
LoGo [6] 0.3969 N/A N/A 0.7406 N/A N/A
ForecastQA [17] Bert N/A 0.3901 N/A N/A 0.7389 N/A
@@
Fine-tune CoH [25] Llama2-7b 0.4856 N/A N/A 0.7763 N/A N/A
GenTKG [24] 0.4785 N/A N/A 0.7907 N/A N/A
@@ 3.1.2 Implementation Details
For the rule-based method, we generate corresponding summaries for all documents in advance with Vicuna-7b. Notably, we further apply fine-tuning (instruction tuning with QLoRA [33]) to these open-source LLMs. Since the retrieved history directly retrieves news text, we utilize gpt-3.5-turbo to generate summaries for the retrieved news text set and keep the retrieval models fixed without tuning. The retrieval models include three different retrievers: BM25 [34], Contriever [35], and LlamaIndex [36].To ensure the fairness of the experiment, for all the baselines and LLM-based methods, we search historical lengths from $\{3,7,15,30,90\}$. Moreover, considering the token length constraints of LLMs, we we limit the number of historical events and texts to 20 and 5, respectively.
3.2 Performance Comparison (RQ1)
@@
3.2.1 Performance w.r.t. Various Inputs and Outputs.
@@
Table 6 presents a comparison of overall performance for both object entity prediction and relation prediction tasks.From these results, we make the following observations. First, the LLM-based method proposed in this paper achieves the highest performance under the fine-tuned setting. These preliminary findings suggest that fine-tuning LLMs holds substantial promise for event forecasting, indicating that more sophisticated fine-tuning strategies warrant further investigation. Compared to the other fine-tuned LLM-based methods such as CoH and GenTKG, the introduction of complex events and news text provides more fine-grained information for temporal event forecasting. Moreover, enhanced with the retrieval modules, LLM-based methods in zero-shot setting can achieve not bad performance for both tasks of tail entity and relation prediction. Nevertheless, they still under-perform conventional non-LLM-based methods on all three formats of inputs, aka. Graph-only, Text-only, and Mixed, demonstrating that non-LLM-based methods remain competitive. Second, surprisingly, incorporating raw text into the input does not demonstrate positive effects in zero-shot setting. The performance on the text-only setting is obviously worse than the graph-only and mixed settings. This may be because the raw textual inputs are longer than the graph-only inputs, resulting in increasing difficulty in reasoning. Third, comparing the two forecasting tasks, the absolute performance of relation prediction is higher than object entity prediction by an obvious margin, indicating that forecasting entities are more challenging than relations in the domain of international cooperation and conflict events. The possible reasons are from two aspects. First, the entity set is much larger than the relation set, therefore, predicting entity is more difficult than relation given a larger candidate set. Second, when predicting relations, the input question includes both the subject and object entity, while the subject entity and relation are the input when predicting object entity. We deem that entities retain more information than relation in event forecasting, thus the question of relation prediction could offer more information and is easier.
3.2.2 Performance w.r.t. Backbone LLMs
@@
The capability of the backbone LLM is critical for robust performance across various tasks. In addition to the commercial proprietary LLMs, we are also interested in the performance of open-source LLMs. These open-source LLMs can be adapted, re-configured, and more importantly, safer than commercial LLMs due to privacy issues. Accordingly, we select two widely adopted open-source LLMs, i.e., Llama2-7b and Vicuna-7b, for evaluation. The results, summarized in Table 7, reveal that both Llama2-7b and Vicuna-7b exhibit significantly lower performance than GPT-3.5-turbo under the zero-shot setting, highlighting the inherent capacity gap between smaller open-source models and more powerful commercial models. Nevertheless, Under the fine-tuning setting, all the open-source models exhibit significant performance gains. They are not only better than the zero-shot version counterparts but also outperform gpt-3.5-turbo by a large margin. This phenomenon indicates that supervised instruction fine-tuning is crucial for enhancing the ability of LLM to learn temporal relational patterns. Additionally, the performance of Vicuna-7b outperforms Llama2-7b, implying that advanced LLM capabilities are more effective at capturing evolutionary patterns in historical events.
@@
Table 7: Performance comparison w.r.t. different backbone LLMs and fine-tuning.
@@
Model Type (rule-based history) Backbone Graph-only Text-only Mixed
Zero-shot Llama2-7b 0.2225 0.2639 0.2157
Vicuna-7b 0.2027 0.2384 0.2524
gpt-3.5-turbo 0.3290 0.3233 0.3154
Fine-tune Llama2-7b 0.5057 0.4892 0.4977
Vicuna-7b 0.5271 0.5266 0.5798
3.2.3 Performance w.r.t. Various Negative Sampling Strategies
Given the large entity set and limited context length of LLMs, we simplify the event forecasting settings as an MCQ problem, i.e., choose the correct answer from six options. However, previous settings, such as TKG works [13, 6], treat the whole entity (relation) set as options during forecasting, aligning well with the real application scenario of the open world. Consequently, our simplified setting of MCQ could inevitably induce bias w.r.t. different sampling strategies for negative (wrong) options generation. To investigate the impact of negative sampling strategy, we introduce two more strategies in addition to the default strategy history, i.e., 1) global, which randomly sample entity from the entire entity set; and 2) generated, which are newly generated by gpt-3.5-turbo and not same with the ground-truth answer  [^7] The generated entities include some entities that are not in the current entity set, therefore, non-LLM methods are not applicable to this setting due to they cannot deal with new entities). Nevertheless, we do not force the generated entities to be totally new, which means the generated entities may be overlapped with existing entities.. The results are shown in the right hand side of Table 6, from which we can observe several interesting findings. First, among all the three strategies, all the methods perform worst on the history setting, showing that our default setting of sampling negative options from event history is the most challenging setting. Second, global demonstrates the highest performance, showing to be the easiest setting. Meanwhile, generated is in the middle, with a balanced level of difficulty. Finally and more interestingly, LLM-based models’ performance on global outperforms all the non-LLM methods. This may be because LLMs are superior in answering general easy questions with common knowledge, while less capable in dealing with hard questions in specific domains.
Table 8: Performance comparison w.r.t. various negative sampling strategies. ”N/A” means ”Not Applicable”.
Backbone Model Negative Sampling Strategy
history global genertaed
GNN ConvTransE 0.3737 0.5237 N/A
RGCN 0.3777 0.5272 N/A
RE-GCN 0.3879 0.5407 N/A
LoGo 0.3969 0.5702 N/A
gpt-3.5-turbo rule-based history 0.3290 0.6189 0.4632
retrieved history 0.3533 0.6433 0.4751
3.3 Effects of RAG (RQ2)